OpenStudy (anonymous):

the length of a rectangular frame is 6cm more than the width. the area inside the frame is 55sq cm. find width of frame.

OpenStudy (anonymous):

It is 5 (:

OpenStudy (anonymous):

haha i think it really is!

OpenStudy (anonymous):

yup, can i have a madal? i gave ya one for 2 plus 2 lol

OpenStudy (anonymous):

2(6+x)+2x=55 12+2x+2x 12+4x=55 12-12+4x=55-12 4x/4=43/4=10.75

OpenStudy (anonymous):

Im pretty sure I got this one right